# Quantum adiabatic optimization with Rydberg arrays: localization phenomena and encoding strategies

### v1

@@ Appendix B Effective Hamiltonian for vertex-wires
Here, we detail the derivation of the effective Hamiltonian for a single vertex-wire [Eq. 4] and for two vertex-wires intersecting via a gadget [Eq.5]. For these setups, it is convenient to rewrite $H_{\rm Rydberg}$ in Eq. (2) in the blockade limit as
@@
We treat ${H_{1}=H_{\Omega}+H_{\epsilon}}$ as small correction of $H_{0}$. In second-order perturbation theory, the dynamics within the lowest energy subspace of $H_{0}$ (of energy $E_{0}$) are described by the effective Hamiltonian
@@
B.1 Vertex-wire
For the vertex-wire in Fig. 1(b), the lowest energy subspace of $H_{0}$ comprises the ${L=N/2+1}$ single-domain-wall states $|j\rangle_{d}$ shown in Fig. 5(a) (with energy ${E_{0}=-N\Delta/2}$), and ${P=\sum_{j=1}^{L=N/2+1}|j\rangle_{d}\langle j|}$. The effective Hamiltonian in Eq. (13) takes the form
@@ B.1 Vertex-wire
Here, $H_{0}$ gives rise to a constant energy term $E_{0}$, while $H_{\epsilon}=\epsilon_{i}n_{1}+\epsilon_{f}n_{N}$ generates both a constant energy term $(\epsilon_{i}+\epsilon_{f})$ and a non-trivial diagonal term ${-(\epsilon_{i}|1\rangle_{d}\langle 1|+\epsilon_{f}|L\rangle_{d}\langle L|)}$. Then, $H_{\Omega}^{2}$ is responsible for two different second-order processes. In the first process, an atom in the Rydberg state is de-excited, while a neighboring atom transitions from the ground to the Rydberg state. This effectively results in a domain-wall hopping of amplitude $t=\Omega^{2}/\Delta$. In the second process, an atom in the Rydberg state is de-excited to the ground state and then re-excited. This lowers the energy by a constant term $-Nt/2$. Thus, Eq. (13) returns Eq. (4) up to a constant diagonal term of $\epsilon_{i}+\epsilon_{f}-\left(\Delta+t\right)N/2$. This effective Hamiltonian holds in the limit of $\left|\epsilon_{i}\right|,\left|\epsilon_{f}\right|,\Omega\ll\Delta$.
B.2 Two intersecting vertex-wires
For two vertex-wires intersecting via a gadget [see Fig. 1(c,d)], the lowest energy eigenstates of $H_{0}$ are characterized by one domain wall per vertex-wire and are depicted in Fig. 5(b,c). The effective Hamiltonian in Eq. (13) takes the form
@@ B.2 Two intersecting vertex-wires
where $Q_{w}$ ($Q_{g}$) is the projector onto the subspace constituted by the states derived from the domain-wall states by de-exciting an excited atom to the ground state along the vertex-wire (in the gadget).
Here, $H_{0}$ generates a constant energy term $E_{0}$, while $H_{\epsilon}=\sum_{w=x,y}\left(\epsilon_{i,w}n_{1,w}+\epsilon_{f,w}n_{N,w}\right)$ — where the subscripts ‘$1,w$’ (‘$N,w$’) label the first (last) atom on the vertex-wire $w$ — generates both a constant energy term $\sum_{w=x,y}(\epsilon_{i,w}+\epsilon_{f,w})$ and a non-trivial diagonal term $-(\sum_{j_{y}}\epsilon_{i,x}|1,j_{y}\rangle_{d}\langle 1,j_{y}|+\epsilon_{f,x}|L_{x},j_{y}\rangle_{d}\langle L_{x},j_{y}|+\sum_{j_{x}}\epsilon_{i,y}|j_{x},1\rangle_{d}\langle j_{x},1|+\epsilon_{f,y}|j_{x},L_{y}\rangle_{d}\langle j_{x},L_{y}|)$. Then, $H_{\Omega}Q_{w}H_{\Omega}$ and $H_{\Omega}Q_{g}H_{\Omega}$ generate the two second-order processes described below.
In the first process, $H_{\Omega}Q_{w}H_{\Omega}$ ($H_{\Omega}Q_{g}H_{\Omega}$) gives rise to the hopping of a single domain wall along one vertex-wire (across the gadget), with amplitude $t=\Omega^{2}/\Delta$ ($t^{\prime}=\Omega^{2}/\Delta_{g}$). Notably, when both domain walls are located at the gadget, $H_{\Omega}Q_{g}H_{\Omega}$ generates a simultaneous hopping of the two domain walls. Indeed, for these states, one of the gadget atoms is excited, while all four vertex-wire atoms attached to the gadget are in the ground state. As a result, $H_{\Omega}Q_{g}H_{\Omega}$ can de-excite the excited gadget atom and excite any other gadget atom, giving rise to the following hopping matrix elements: $|\zeta_{x},\zeta_{y}\rangle_{d}\langle\zeta_{x}+1,\zeta_{y}+1|$ and $|\zeta_{x}+1,\zeta_{y}\rangle_{d}\langle\zeta_{x},\zeta_{y}+1|$ for the crossing gadget, and only $|\zeta_{x}+1,\zeta_{y}\rangle_{d}\langle\zeta_{x},\zeta_{y}+1|$ for the CWE gadget.
@@
Appendix C Vertex-wire and continuum model
Here, we discuss the eigenvalue problem for the effective Hamiltonian of a vertex-wire in Eq. (4) and show that, for $\epsilon_{i}=\epsilon_{f}=0$ and $\epsilon_{i}=\epsilon_{f}=t$, it can be mapped to a 1D-Helmholtz equation with homogeneous Dirichlet boundary conditions (DBC) and homogeneous Neumann boundary conditions (NBC), respectively.
@@ C.1.1 Solutions to the eigenvalue problem
and $m=0,\dots,L-1$. Boundary conditions strongly affect the ground state of the system. For $r_{i}=r_{f}=0$, the ground-state population is minimum at the extremes and maximum in the middle of the vertex-wire. In contrast, for $r_{i}=r_{f}=1$, the ground-state population is constant along the entire system (namely, the ground state is the equal-weight superposition state).
@@
For $r_{i}$ or(and) $r_{f}$ larger than $1$, Eq. (16) admits one(two) imaginary $k=i\kappa$ solutions. In the $\kappa L\to\infty$ limit, we find: $\kappa_{i}=\log{r_{i}}$ when $r_{i}>1$, and $\kappa_{f}=\log{r_{f}}$ when $r_{f}>1$. The associated wavefunctions are
@@ Appendix D Further analyses on the crossing setup
D.1 Effective 1D model: single vertex-wire with two modified detuning
Here, we develop an effective theory for a vertex-wire of ${N+2}$ atoms with inhomogeneous detuning: $\Delta_{g}$ on the two atoms at the middle of the vertex-wire and $\Delta$ elsewhere. The two midpoint atoms mimic a gadget and divide the vertex-wire into two legs, each with an even number of atoms $m=N/2$. This model effectively describes the dynamics of a single domain wall along a vertex-wire in the crossing setup (when the position of the domain wall along the other vertex-wire is considered to be fixed).
@@ D.1 Effective 1D model: single vertex-wire with two modified detuning
For $\Delta_{g}>0$, the lowest energy subspace of $H_{0}$ comprises single-domain-wall states of the vertex-wire, and Eq. (13) returns the following effective Hamiltonian (up to a constant diagonal term $-\left[\left(\Delta+t\right)N/2+\Delta_{g}+t^{\prime}\right]\mathbb{I}$):
@@
with $j=1,\dots,L$ (where $L=N/2+2$), $\zeta=N/2$, $t=\Omega^{2}/\Delta$, and $t^{\prime}=\Omega^{2}/\Delta_{g}$. This effective Hamiltonian is the same as the one of the crossing setup, represented in Fig. 15(a), when fixing the position of one of the two domain walls. When $0<\Delta_{g}<\Delta$, the ground state is localized at $\zeta$ and $\zeta+1$ (around the gadget) and exponentially decays in the two legs. In particular, for an infinite vertex-wire, the ground state has energy $-(t^{\prime 2}+t^{2})/t^{\prime}$, and it is given by (up to a normalization factor):
@@ D.2 Phase diagram
Figure 16: (a) Finite-size scaling analysis of the minimum energy gap of the crossing setup along the standard protocol. We consider two vertex-wires of $N$ atoms, with $N$ ranging from $100$ (light orange) to $172$ (dark orange), intersecting at the middle ($m_{x}=m_{y}=N/2$), and $\Delta_{g}=3\Delta$. The data collapse to Eq. (32) for $z\approx 0.91$, $\nu\approx 1.07$, and $\lambda_{c}\approx 1.32$. (b) Ground state energy, up to a constant $(\Delta+t)N$, where $t=\Omega^{2}/\Delta$, as a function of $\Delta_{g}/\Delta$ for $N=50$ and $\Delta=10\Omega$ obtained using DMRG (blue dots) and the effective theory in Eq. (29) (orange crosses). The black lines are the analytical predictions $E_{\Delta_{g}<0}=-8t$ and $E_{\Delta_{g}>0}=-\left(\Delta_{g}+\sqrt{\Delta_{g}^{2}+16\Omega^{2}}\right)/2$. The values in the plot are shown for $\Omega=1$. (c) Rydberg excitation density in the gadget $n_{\rm gad}$ as a function of $\Delta/\Omega$ and $\Delta_{g}/\Omega$. (d) Entanglement entropy $S$ between one leg and the rest of the system as a function of $\Delta/\Omega$ and $\Delta_{g}/\Omega$.
@@
Here, $s_{i}$ is the singular value at the bond that connects one leg to the gadget tensor. We show $n_{\mathrm{gad}}$ in Fig. 16(c) and $S$ in Fig. 16(d) as a function of $\Delta/\Omega$ and $\Delta_{g}/\Omega$. In the following, we provide an interpretation of these quantities based on the effective 1D model developed in the previous subsection for a single vertex-wire with two modified detuning.
For $\Delta_{g}<0$, the gadget is empty ($n_{\rm gad}\approx 0$). When going from the disordered to the ordered phase, each leg is decoupled from the rest of the system (since $S\approx 0$) and independently undergoes the $\mathbb{Z}_{2}$ spontaneous symmetry breaking of a one-dimensional Rydberg chain. Consequently, the system enters an ordered regime characterized by a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Within this ordered regime, a perturbative description is then possible when $\Omega\ll\Delta,\left|\Delta_{g}\right|$. The effective Hamiltonian is obtained by generalizing the single-vertex-wire Hamiltonian in Eq. (28) to two vertex-wires. In this regime, the ground-state energy is given by $E_{\Delta_{g}<0}=-8t$ (twice that of the single vertex-wire).
At the opposite limit $\Delta_{g}\gtrsim\Delta$, the gadget hosts a Rydberg excitation ($n_{\rm gad}\approx 1$). This has the effect of coupling opposite legs, which behave as a single vertex-wire and undergo (in pair) the $\mathbb{Z}_{2}$ symmetry breaking. Consequently, at $\Delta/\Omega\sim 1.3$, the system enters an ordered regime with a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Here, the entanglement entropy between a single leg and the rest of the system (specifically, with the opposite leg) is maximum. In the $\Delta_{g}\gg\Delta$ limit, a perturbative treatment of the system generalizes the results of the single-wire analysis [see Eq. (26)] and predicts a ground-state energy equal to $E_{\Delta_{g}\gg\Delta}=-4t$ (twice that of the single vertex-wire).
The active and inactive regimes are separated by an intermediate region ($0<\Delta_{g}<\Delta$). Here, the occupation of the gadget varies around $n_{\rm gad}\approx 0.5$ [see Fig. 16(c)], and the Rydberg excitation is equally superpositioned among the four states depicted in the inset of Fig. 15(a). In the $\left|\Delta_{g}\right|\gg\Omega$ limit, the system can be studied perturbatively, generalizing the single-wire Hamiltonian in Eq. (29) to two vertex-wires. In particular, the simplified three-level model is now adapted to a five-level one — one empty gadget state plus the four states in Fig. 15(a) — which predicts ground-state energy equal to $E_{\Delta_{g}>0}=-\left(\Delta_{g}+\sqrt{\Delta_{g}^{2}+16\Omega^{2}}\right)/2$.
@@ D.3 Ising transition
In Sec. III.1, we showed that the minimum energy gap, at the transition between the disordered phase and the active regime ($\mathbb{Z}_{2}\times\mathbb{Z}_{2}$), scales as $\sim L^{-z}$ with $z\approx 1$. To gain a better understanding of this phase transition, we perform a finite-size scaling analysis [88, 89]. We use the following ansatz for the minimum gap:

### v2

@@ Appendix B Effective Hamiltonian for vertex wires
Here, we detail the derivation of the effective Hamiltonian for a single vertex wire [Eq. 4] and for two vertex wires intersecting via a gadget [Eq.5]. For these setups, it is convenient to rewrite $H_{\rm Rydberg}$ in Eq. (2) in the blockade limit as
@@
We treat ${H_{1}=H_{\Omega}+H_{\epsilon}}$ as small correction of $H_{0}$. In second-order perturbation theory, the dynamics within the lowest-energy subspace of $H_{0}$ (of energy $E_{0}$) are described by the effective Hamiltonian
@@
B.1 Vertex wire
For the vertex wire in Fig. 1(b), the lowest energy subspace of $H_{0}$ comprises the ${L=N/2+1}$ single-domain-wall states $|j\rangle_{d}$ shown in Fig. 5(a) (with energy ${E_{0}=-N\Delta/2}$), and ${P=\sum_{j=1}^{L=N/2+1}|j\rangle_{d}\langle j|}$. The effective Hamiltonian in Eq. (13) takes the form
@@ B.1 Vertex wire
Here, $H_{0}$ gives rise to a constant energy term $E_{0}$, while $H_{\epsilon}=\epsilon_{i}n_{1}+\epsilon_{f}n_{N}$ generates both a constant energy term $(\epsilon_{i}+\epsilon_{f})$ and a nontrivial diagonal term ${-(\epsilon_{i}|1\rangle_{d}\langle 1|+\epsilon_{f}|L\rangle_{d}\langle L|)}$. Then, $H_{\Omega}^{2}$ is responsible for two different second-order processes. In the first process, an atom in the Rydberg state is de-excited, while a neighboring atom transitions from the ground to the Rydberg state. This effectively results in a domain-wall hopping of amplitude $t=\Omega^{2}/\Delta$. In the second process, an atom in the Rydberg state is de-excited to the ground state and then re-excited. This lowers the energy by a constant term $-Nt/2$. Thus, Eq. (13) returns Eq. (4) up to a constant diagonal term of $\epsilon_{i}+\epsilon_{f}-\left(\Delta+t\right)N/2$. This effective Hamiltonian holds in the limit of $\left|\epsilon_{i}\right|,\left|\epsilon_{f}\right|,\Omega\ll\Delta$.
B.2 Two intersecting vertex wires
For two vertex wires intersecting via a gadget [see Fig. 1(c,d)], the lowest-energy eigenstates of $H_{0}$ are characterized by one domain wall per vertex wire and are depicted in Fig. 5(b,c). The effective Hamiltonian in Eq. (13) takes the form
@@ B.2 Two intersecting vertex wires
where $Q_{w}$ ($Q_{g}$) is the projector onto the subspace constituted by the states derived from the domain-wall states by de-exciting an excited atom to the ground state along the vertex wire (in the gadget).
Here, $H_{0}$ generates a constant energy term $E_{0}$, while $H_{\epsilon}=\sum_{w=x,y}\left(\epsilon_{i,w}n_{1,w}+\epsilon_{f,w}n_{N,w}\right)$ — where the subscripts ‘$1,w$’ (‘$N,w$’) label the first (last) atom on the vertex wire $w$ — generates both a constant energy term $\sum_{w=x,y}(\epsilon_{i,w}+\epsilon_{f,w})$ and a nontrivial diagonal term $-(\sum_{j_{y}}\epsilon_{i,x}|1,j_{y}\rangle_{d}\langle 1,j_{y}|+\epsilon_{f,x}|L_{x},j_{y}\rangle_{d}\langle L_{x},j_{y}|+\sum_{j_{x}}\epsilon_{i,y}|j_{x},1\rangle_{d}\langle j_{x},1|+\epsilon_{f,y}|j_{x},L_{y}\rangle_{d}\langle j_{x},L_{y}|)$. Then, $H_{\Omega}Q_{w}H_{\Omega}$ and $H_{\Omega}Q_{g}H_{\Omega}$ generate the two second-order processes described below.
In the first process, $H_{\Omega}Q_{w}H_{\Omega}$ ($H_{\Omega}Q_{g}H_{\Omega}$) gives rise to the hopping of a single domain wall along one vertex wire (across the gadget), with amplitude $t=\Omega^{2}/\Delta$ ($t^{\prime}=\Omega^{2}/\Delta_{g}$). Notably, when both domain walls are located at the gadget, $H_{\Omega}Q_{g}H_{\Omega}$ generates a simultaneous hopping of the two domain walls. Indeed, for these states, one of the gadget atoms is excited, while all four vertex-wire atoms attached to the gadget are in the ground state. As a result, $H_{\Omega}Q_{g}H_{\Omega}$ can de-excite the excited gadget atom and excite any other gadget atom, giving rise to the following hopping matrix elements: $|\zeta_{x},\zeta_{y}\rangle_{d}\langle\zeta_{x}+1,\zeta_{y}+1|$ and $|\zeta_{x}+1,\zeta_{y}\rangle_{d}\langle\zeta_{x},\zeta_{y}+1|$ for the crossing gadget, and only $|\zeta_{x}+1,\zeta_{y}\rangle_{d}\langle\zeta_{x},\zeta_{y}+1|$ for the CWE gadget.
@@
Appendix C vertex wire and continuum model
Here, we discuss the eigenvalue problem for the effective Hamiltonian of a vertex wire in Eq. (4) and show that, for $\epsilon_{i}=\epsilon_{f}=0$ and $\epsilon_{i}=\epsilon_{f}=t$, it can be mapped to a 1D-Helmholtz equation with homogeneous Dirichlet boundary conditions (DBCs) and homogeneous Neumann boundary conditions (NBCs), respectively.
@@ C.1.1 Solutions to the eigenvalue problem
and $m=0,\dots,L-1$. Boundary conditions strongly affect the ground state of the system. For $r_{i}=r_{f}=0$, the ground-state population is minimum at the extremes and maximum in the middle of the vertex wire. In contrast, for $r_{i}=r_{f}=1$, the ground-state population is constant along the entire system (namely, the ground state is the equal-weight superposition state).
@@
For $r_{i}$ or(and) $r_{f}$ larger than $1$, Eq. (16) admits one(two) imaginary $k=i\kappa$ solutions. In the $\kappa L\to\infty$ limit, we find $\kappa_{i}=\log{r_{i}}$ when $r_{i}>1$, and $\kappa_{f}=\log{r_{f}}$ when $r_{f}>1$. The associated wavefunctions are
@@ Appendix D Further analyses on the crossing setup
D.1 Effective 1D model: single vertex wire with two modified detuning
Here, we develop an effective theory for a vertex wire of ${N+2}$ atoms with inhomogeneous detuning: $\Delta_{g}$ on the two atoms at the middle of the vertex wire and $\Delta$ elsewhere. The two midpoint atoms mimic a gadget and divide the vertex wire into two legs, each with an even number of atoms $m=N/2$. This model effectively describes the dynamics of a single domain wall along a vertex wire in the crossing setup (when the position of the domain wall along the other vertex wire is considered to be fixed).
@@ D.1 Effective 1D model: single vertex wire with two modified detuning
For $\Delta_{g}>0$, the lowest-energy subspace of $H_{0}$ comprises single-domain-wall states of the vertex wire, and Eq. (13) returns the following effective Hamiltonian (up to a constant diagonal term $-\left[\left(\Delta+t\right)N/2+\Delta_{g}+t^{\prime}\right]\mathbb{I}$):
@@
with $j=1,\dots,L$ (where $L=N/2+2$), $\zeta=N/2$, $t=\Omega^{2}/\Delta$, and $t^{\prime}=\Omega^{2}/\Delta_{g}$. This effective Hamiltonian is the same as the one of the crossing setup, represented in Fig. 15(a), when fixing the position of one of the two domain walls. When $0<\Delta_{g}<\Delta$, the ground state is localized at $\zeta$ and $\zeta+1$ (around the gadget) and exponentially decays in the two legs. In particular, for an infinite vertex wire, the ground state has energy $-(t^{\prime 2}+t^{2})/t^{\prime}$, and it is given by (up to a normalization factor):
@@ D.2 Phase diagram
Figure 16: (a) Finite-size scaling analysis of the minimum energy gap of the crossing setup along the standard protocol. We consider two vertex wires of $N$ atoms, with $N$ ranging from $100$ (light orange) to $172$ (dark orange), intersecting at the middle ($m_{x}=m_{y}=N/2$), and $\Delta_{g}=3\Delta$. The data collapse to Eq. (32) for $z\approx 0.91$, $\nu\approx 1.07$, and $\lambda_{c}\approx 1.32$. (b) Ground-state energy, up to a constant $(\Delta+t)N$, where $t=\Omega^{2}/\Delta$, as a function of $\Delta_{g}/\Delta$ for $N=50$ and $\Delta=10\Omega$ obtained using DMRG (blue dots) and the effective theory in Eq. (29) (orange crosses). The black lines are the analytical predictions $E_{\Delta_{g}<0}=-8t$ and $E_{\Delta_{g}>0}=-\left(\Delta_{g}+\sqrt{\Delta_{g}^{2}+16\Omega^{2}}\right)/2$. The values in the plot are shown for $\Omega=1$. (c) Rydberg excitation density in the gadget $n_{\rm gad}$ as a function of $\Delta/\Omega$ and $\Delta_{g}/\Omega$. (d) Entanglement entropy $S$ between one leg and the rest of the system as a function of $\Delta/\Omega$ and $\Delta_{g}/\Omega$.
@@
Here, $s_{i}$ is the singular value at the bond that connects one leg to the gadget tensor. We show $n_{\mathrm{gad}}$ in Fig. 16(c) and $S$ in Fig. 16(d) as a function of $\Delta/\Omega$ and $\Delta_{g}/\Omega$. In the following, we provide an interpretation of these quantities based on the effective 1D model developed in the previous subsection for a single-vertex wire with two modified detuning.
For $\Delta_{g}<0$, the gadget is empty ($n_{\rm gad}\approx 0$). When going from the disordered to the ordered phase, each leg is decoupled from the rest of the system (since $S\approx 0$) and independently undergoes the $\mathbb{Z}_{2}$ spontaneous symmetry breaking of a one-dimensional Rydberg chain. Consequently, the system enters an ordered regime characterized by a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Within this ordered regime, a perturbative description is then possible when $\Omega\ll\Delta,\left|\Delta_{g}\right|$. The effective Hamiltonian is obtained by generalizing the single-vertex wire Hamiltonian in Eq. (28) to two vertex wires. In this regime, the ground-state energy is given by $E_{\Delta_{g}<0}=-8t$ (twice that of the single vertex wire).
At the opposite limit $\Delta_{g}\gtrsim\Delta$, the gadget hosts a Rydberg excitation ($n_{\rm gad}\approx 1$). This has the effect of coupling opposite legs, which behave as a single vertex wire and undergo (in pair) the $\mathbb{Z}_{2}$ symmetry breaking. Consequently, at $\Delta/\Omega\sim 1.3$, the system enters an ordered regime with a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. Here, the entanglement entropy between a single leg and the rest of the system (specifically, with the opposite leg) is maximum. In the $\Delta_{g}\gg\Delta$ limit, a perturbative treatment of the system generalizes the results of the single-wire analysis [see Eq. (26)] and predicts a ground-state energy equal to $E_{\Delta_{g}\gg\Delta}=-4t$ (twice that of the single vertex wire).
The active and inactive regimes are separated by an intermediate region ($0<\Delta_{g}<\Delta$). Here, the occupation of the gadget varies around $n_{\rm gad}\approx 0.5$ [see Fig. 16(c)], and the Rydberg excitation is equally superpositioned among the four states depicted in the inset of Fig. 15(a). In the $\left|\Delta_{g}\right|\gg\Omega$ limit, the system can be studied perturbatively, generalizing the single-wire Hamiltonian in Eq. (29) to two vertex wires. In particular, the simplified three-level model is now adapted to a five-level one — one empty gadget state plus the four states in Fig. 15(a) — which predicts ground-state energy equal to $E_{\Delta_{g}>0}=-\left(\Delta_{g}+\sqrt{\Delta_{g}^{2}+16\Omega^{2}}\right)/2$.
@@ D.3 Ising transition
In Sec. III.1, we showed that the minimum energy gap, at the transition between the disordered phase and the active regime ($\mathbb{Z}_{2}\times\mathbb{Z}_{2}$), scales as $L^{-z}$ with $z\approx 1$. To gain a better understanding of this phase transition, we perform a finite-size scaling analysis [94, 95]. We use the following ansatz for the minimum gap: